# From Infants to AI: Incorporating Infant-like Learning in Models Boosts Efficiency and Generalization in Learning Social Prediction Tasks

### v2

@@ Abstract
Early in development, infants learn a range of useful concepts, which can be challenging from a computational standpoint. This early learning comes together with an initial understanding of aspects of the meaning of concepts, e.g., their implications, causality, and using them to predict likely future events. All this is accomplished in many cases with little or no supervision, and from relatively few examples, compared with current network models. In learning about objects and human-object interactions, early acquired and possibly innate concepts are often used in the process of learning additional, more complex concepts. In the current work, we model how early-acquired concepts are used in the learning of subsequent concepts, and compare the results with standard deep network modeling. We focused in particular on the use of the concepts of animacy and goal attribution in learning to predict future events. We show that the use of early concepts in the learning of new concepts leads to better learning (higher accuracy) and more efficient learning (requiring less data). We further show that this integration of early and new concepts shapes the representation of the concepts acquired by the model. The results show that when the concepts were learned in a human-like manner, the emerging representation was more useful, as measured in terms of generalization to novel data and tasks. On a more general level, the results suggest that there are likely to be basic differences in the conceptual structures acquired by current network models compared to human learning.
@@
Keywords: Computational Modeling; Infant Learning; Artificial Intelligence; Social Cognition; Machine Learning
Introduction
@@
Already in early development, infants learn a broad range of useful concepts and visual tasks, which can be challenging from a computational standpoint. For example, in the domains of objects and human-object interactions (on which we focus), infants learn during the first year of life to recognize hands, their configuration, and their interactions with objects (?, ?, ?, ?), a task where significant and meaningful features can be non-salient and highly variable and therefore difficult to learn.
@@
Infants learn, in an unsupervised manner, to perform figure-ground segmentation, which took years and a large effort to develop (?, ?). In the domain of dealing with other people, starting at 3-6 months of age, infants learn to detect and follow another person’s gaze and establish joint attention, on the basis of head orientation, and later, eye direction (?, ?, ?, ?). This task is difficult, because ‘gaze’ is not a physical entity that appears explicitly in the image, and cues for gaze direction can be subtle and difficult to extract and use.
@@
The learning of early concepts comes together with an initial understanding of aspects of the concepts’ meaning, in terms of their implications, causality, or use for predicting likely future events. For example, in learning to identify hands, infants also learn that hands cause other objects to move and change location (?, ?, ?). All this learning is obtained in many cases with little or no supervision, and from relatively few examples, compared with current models (?, ?). In learning about objects and human-object interactions, early learned concepts are often used in the process of acquiring additional related concepts. For example, the relation of in front/behind appears to be a prerequisite to the subsequent learning of containment (?, ?). Learning to identify the direction of gaze is used to predict future actions (?, ?), and it later plays a role in the development of communication and language (?, ?).
@@
In the current work, we model how early-acquired concepts are used in the learning of subsequent concepts. In particular, we focus on using the concepts of animacy and goal attribution in learning to predict future events. The main approach is to compare the learning of new concepts in two similar networks, where only one of the two uses a decomposition, similar to what was shown in infant studies, of learning a visual task by learning a simple concept first and then using it for learning the new task. For example, we compare the representation of the concept ‘animate’ and how it is used by the model, in the two different learning schemes: learning ‘animate’ first, and then using it to predict future behavior, compared to combined learning of animacy together with predicting future behavior.
@@
We refer to the version that uses human-like decomposition of concepts as the ‘cognitive’ model, and the standard, end-to-end training as the ‘naive’ model. We compare the cognitive and standard approaches along two main directions: one is the performance of the trained models, and the other is a comparison of what was learned and represented in the two types of models. Briefly, in terms of model performance, the results show that the use of early concepts in the learning of new concepts leads to better learning (higher accuracy) and more efficient learning (requiring less data). In terms of the learned concept representations, the results show that when the concepts are learned in a human-like manner, the emerging representation is more useful, as measured in terms of generalization to novel data and tasks. In the final discussion, we examine the possible source of the differences (decomposition and early concepts) and potential implications of the findings to the training of artificial models.
The main contributions of this work are as follows:
@@
  • Our study is the first (as far as we know) to test the potential computational advantage of a basic characteristic of infants’ learning in terms of systematically using a range of early-acquired concepts in the learning of subsequent, more complex ones.
  • We show that the infant-like learning model (in the sense above) reaches better performance on a test task in terms of accuracy as well as efficiency (size of the training set).
  • We show evidence that the use of early-learned concepts in the learning of more complex ones plays a role in shaping the final representations produced during the learning process.
Related Work
Infants’ Understanding of Animacy and Goal Attribution
@@
An early and influential contribution to understanding infants’ goal encoding is Woodward’s work (?, ?), which demonstrated that infants attribute goals to agents or animate actors, but not to inanimate objects. This finding has been validated and elaborated through other experiments and variations (?, ?, ?, ?). A study by Luo and Baillargeon (?, ?) extended the understanding of goal attribution by showing that infants can attribute goals to self-propelled objects. Further research illustrated that infants can generalize goal attribution to novel actions when provided with sufficient contextual cues (?, ?). Additionally, infants exhibit a significant understanding of objects and their properties from an early age (?, ?, ?, ?), expecting objects to follow physical rules (?, ?, ?).
Related Computational Models of Infants’ Intuitive Psychology
@@
A recent study (?, ?) investigated goal attribution in infants, both empirically and in models. The study found that infants anticipate agents’ actions to be directed at objects rather than locations. In contrast, AI models often target locations, highlighting the need to integrate infants’ understanding into models to better replicate human behavior. Li’s work (?, ?) introduced additional tasks related to agents and objects, along with a self-supervised model. This work demonstrated limitations of current neural network models in goal attribution tasks.
@@
Several works (?, ?, ?, ?) have addressed infant-like tasks, using the Baby Intuitions Benchmark (BIB) (?, ?), a dataset specifically designed for evaluating developmental tasks related to agents. These studies employed various approaches, such as transformer-based architectures and Bayesian models, to tackle these tasks.
@@
Goal-Directed Dataset
We designed a dataset inspired by Woodward’s experiments (?, ?), extending it to include tasks beyond those studied in prior research. The dataset consists of sequences of simple scenes (Figure 1), created using icons from ’flaticon.com’. Each frame contains three entities: two inanimate objects and an actor, which can be either animate (e.g., hands, animals, people) or inanimate (e.g., books, flowers).
@@
First Step - Scene Representations
The objective of the first step is to create scene representations for the image sequences. To achieve this, we fine-tuned the BLIP model (?, ?) to generate representations for each frame, encoding the class and location of the three entities (upper-left, upper-right, bottom). For example, the representation for the first frame in the top row of Figure 1 would be: ’cake’, ’bag’, ’girl’. The meaning of the words (e.g., ’cake’, ’bag’, ’girl’) is not used by the model. Instead, the words only serve as placeholders to encode the entities’ identities.
@@
Second Step - Prediction Generation
The prediction task is framed as a classification problem using the BERT model (?, ?). The input to the model is the full sequence representation of the frames, and the model’s output is a prediction of the actor’s future location (’left’ or ’right’) in the next step of the sequence. The results in the following sections evaluate the model’s accuracy in this prediction.
We used large models (BLIP, BERT) in our cognitive model to reflect the complexity of the infant brain—infants, too, are born with a sophisticated neural architecture. We fine-tuned all layers in both models for our specific tasks. To ensure that the models’ pre-trained linguistic knowledge did not drive the observed effects, we also ran a version of the experiment using binary representations (e.g., 0/1 vectors) instead of natural language and obtained similar results. Furthermore, we confirmed that their prior knowledge did not bias the results: the models’ initial performance on our tasks was not significantly different from chance level.
Figure 2: Two-step process of the cognitive and naive models: First, scene representations are created, and then predictions are made. Concepts specific only to the cognitive model are in bold and purple.
Experiments
Experiment 1: Prediction
@@
The primary question was how the two models compare in learning the prediction task. Two important points regarding this comparison are: 1. The naive model could theoretically learn to distinguish between the two types of actors based on the data it observes, as the actor type is consistently correlated with its behavior in both training and test scenarios. 2. Note that the addition, such as the labels ‘animate’ and ‘non-animate’, is provided in a ’bare’ form, without additional context about its implications or how it relates to the model’s predictions. For the model, these labels are arbitrary markers (just as e.g. ’goo’ vs. ’not goo’).
Data
The models were trained using sequences of three frames (Figure 1), employing various combinations of icons, with half of the sequences showing animate actors and half showing non-animate actors. Two dataset sizes were used: a small dataset (320 training examples and 80 test examples) and a large dataset (1280 training, 320 test examples). The test data used the same actors as the training data but introduced novel target objects to evaluate generalization. Both the cognitive and naive networks were trained for 2000 epochs. Each experiment was repeated 15 times with randomly generated sequences, and the results were averaged.
Results
The averaged results of Experiment 1 for both models are shown in Figure 3, comparing their performance on the small and large datasets. Only the first 1000 epochs are shown, as the results stabilized within this range. The cognitive network achieved perfect accuracy after 100 epochs with the smaller dataset and converged to perfect accuracy after 50 epochs with the larger dataset. In contrast, the naive network achieved approximately 80% accuracy after 1000 epochs with the smaller dataset and 87% accuracy with the larger dataset.
@@ Results
Experiment 2: Generalization
Generalization to new tasks and domains is a crucial aspect and a useful measure of intelligent behavior (?, ?, ?). In the previous experiment, the test data used the same actors as in training but introduced new target objects. Here, we extended this to assess the networks’ ability to generalize by evaluating their performance with entirely new actors.
To address generalization, a 5-frame paradigm was introduced (Figure 4). This paradigm first presents the actor’s behavior in the initial three frames and then introduces the same actor with new target objects, requiring the model to predict the actor’s behavior in the next step. Unlike Experiment 1, this setup includes information about the actor’s behavior within the data itself, theoretically enabling both networks to predict an actor’s behavior from a single example. As a result, it is well suited for testing generalization to new actors.
Data
The 5-frame paradigm incorporates the second and third frames from Experiment 1 (Prediction task), along with the frame that features the next step of the sequence, which was not shown in Experiment 1. The first three frames of Experiment 2 already show the actor’s preference. This sequence is extended by adding two new frames that present the same actor with new target objects. The prediction task focuses on the actor’s behavior in relation to these new targets. The models were trained using sequences of five frames, each featuring various combinations of icons, with half of the sequences containing animate actors and the other half containing non-animate actors.
Two task types were generated: Task 1 (T1): A simpler task where the test data use the same actors as in the training data but introduce new target objects. Task 2 (T2): A more challenging task because the test data include entirely new actors and new target objects.
The networks were first trained on T1 and then further fine-tuned under one of two conditions: T1-T1: The same simple task (T1) with additional data. This condition served as a control to assess the effect of additional data without introducing new challenges. T1-T2: The more challenging generalization task (T2). Each task used 640 training examples and 160 test examples. Both networks were trained for 1000 epochs. Each experiment was repeated 15 times with randomly generated sequences, and the results were averaged.
@@
Results
The results for both models are shown in Figure 5, comparing the performance under the following conditions: 1. Learning the initial task (T1). 2. Retraining on the same task with additional data (T1-T1). 3. Generalization to a new task (T1-T2). Only the first 300 epochs are shown, as there was no significant change afterward. The cognitive network outperformed the naive network in both accuracy and learning speed across all tasks: T1-T1 Condition: The cognitive network achieved near-perfect accuracy in the first run and 100% accuracy almost immediately in the second run. The naive network stabilized at 58% accuracy after 1000 epochs in the first run and reached 75% accuracy after 200 epochs in the second run. T1-T2 Condition: For the challenging generalization task, the cognitive network achieved perfect accuracy after just a few epochs. The naive network, however, reached only  65% accuracy, performing worse than in the T1-T1 condition.
@@
Figure 5: Experiment 2: Generalization Results. Average test accuracy of the naive and cognitive networks on Task 1 (T1), retraining on Task 1 (T1-T1), and retraining on Task 2 (T1-T2), with SEM included.
Experiment 3: Decomposition
We address the following question: How is the additional information provided to the cognitive model learned? In earlier experiments, the actor’s type (animate/inanimate) was explicitly provided to the cognitive network. In the current experiment, however, the same information is available to both models. Inspired by evidence that self-propelled motion is a strong indicator of animacy in infants (?, ?), the experiment introduced additional frames to enable the learning of this concept.

### v3

# Concepts Learned Visually by Infants Can Contribute to Visual Learning and Understanding in AI Models

@@ Abstract
Early in development, infants learn to extract surprisingly complex aspects of visual scenes. This early learning comes together with an initial understanding of the extracted concepts, such as their implications, causality, and using them to predict likely future events. In many cases, this learning is obtained with little or no supervision, and from relatively few examples, compared to current network models. Empirical studies of visual perception in early development have shown that in the domain of objects and human-object interactions, early-acquired concepts are often used in the process of learning additional, more complex concepts. In the current work, we model how early-acquired concepts are used in the learning of subsequent concepts, and compare the results with standard deep network modeling. We focused in particular on the use of the concepts of animacy and goal attribution in learning to predict future events in dynamic visual scenes. We show that the use of early concepts in the learning of new concepts leads to better learning (higher accuracy) and more efficient learning (requiring less data), and that the combination of early and new concepts shapes the representation of the concepts acquired by the model and improves its generalization. We further compare advanced vision-language models to a human study in a task that requires an understanding of the behavior of animate vs. inanimate agents, with results supporting the contribution of early concepts to visual understanding. We finally briefly discuss the possible benefits of incorporating aspects of human-like visual learning into computer vision models.
keywords:
@@
Infant Visual Learning; Infant Social Learning; Computational Cognitive Modeling; Computer Vision; Vision-Language Models; Artificial Intelligence
1 Introduction
In early development, infants learn a broad range of useful concepts and visual tasks, which can be challenging from a computational point of view. Our goal in this work is to model aspects of this intriguing capacity and to compare human vision to vision-language models (VLMs) (We use the term VLMs because our focus is on vision-related aspects; however, our discussion also applies to so-called large multimodal models (LMMs).) that do not have this capacity. Finally, we consider the incorporation of this early human vision capacity in future large models.
@@ 1 Introduction
In the domains of objects and human-object interactions (on which we focus), infants learn during the first year of life to recognize hands, their configuration, and their interactions with objects [42, 13, 32], a task where significant and meaningful features can be non-salient and highly variable and therefore difficult to learn.
@@
Infants learn, in an unsupervised manner, to perform figure-ground segmentation, which took years and a great effort to develop [19]. In the domain of dealing with other people, starting at 3-6 months of age, infants learn to detect and follow the gaze of another person and establish joint attention, based on head orientation and later on eye direction [34, 9, 5]. This task is difficult because ‘gaze’ is not a physical entity that appears explicitly in the image, and cues for gaze direction can be subtle and difficult to extract and use.
@@
The learning of early visual concepts comes together with an initial understanding of aspects of the meaning of concepts, in terms of their implications, causality, or use to predict probable future events. For example, in learning to identify hands, infants also learn that hands cause other objects to move and change location [32, 18]. All of this learning is obtained in many cases with little or no supervision, and from relatively few examples, compared to current models [40]. In learning about objects and human-object interactions, early learned concepts are often used in the process of acquiring additional related concepts. For example, the relation of in front/behind appears to be a prerequisite to subsequent learning of containment [40]. Learning to identify the direction of gaze is used to predict future actions [8], and it later plays a role in the development of communication and language [39].
@@
In the current work, we model how early-acquired visual concepts are used in the learning of subsequent concepts. In particular, we focus on using the concepts of animacy and goal attribution in learning to predict future events in dynamic scenes. The main approach is to compare the learning of new concepts in two similar networks, where only one of the two uses a decomposition, similar to what was shown in infant studies, of learning a visual task by learning a simple concept first and then using it for learning the new task. For example, we compare the representation of the concept ‘animate’ and how it is used by the model, in the two different learning schemes: learning ‘animate’ first and then using it to predict future behavior, compared to combined learning of animacy together with prediction of future behavior.
@@
We refer to the version that uses human-like decomposition of concepts as the ‘cognitive’ model, and the standard, end-to-end training, as the ‘naive’ model. We compare the cognitive and standard approaches along two main directions: one is the performance of the trained models, and the other is a comparison of what was learned and represented in the two types of model. Briefly, in terms of model performance, the results show that the use of early concepts in the learning of new concepts leads to better learning (higher accuracy) and more efficient learning (requiring less data). In terms of the learned concept representations, the results show that when the concepts are learned in a human-like manner, the emerging representation is more useful, as measured in terms of generalization to novel data and tasks. We next examine the performance of a range of current large models in tasks that involve the use of animacy and goal attribution in predicting future human behavior. Using a novel human test, we identify limitations of current models in this task. In a final discussion, we examine how early-emerging concepts, incorporated in a hierarchy of concepts, may contribute to better learning. We propose two sources to this contribution – the discovery of useful fundamental concepts, and the role of early concepts in the decomposition of the learning task into simpler components. Finally, we consider the potential benefits of incorporating early human conceptual structure into large vision-language models. The main contributions of this work are as follows:
@@
  • Our study shows the computational advantage of systematically using a range of early acquired visual concepts in the learning of subsequent, more complex ones.
  • We show that our human-like learning model contributes to the performance of standard models in terms of accuracy, efficiency and generalization.
  • Using a new human experiment, we show limitations of current vision-language models in predicting human behaviour in a visual task.
  • Based on our results, we propose a specific way of improving vision-language models by training their visual component on early human visual concepts.
2 Related Work
2.1 Infants’ Understanding of Animacy and Goal Attribution
@@
An early and influential contribution to understanding infants’ goal encoding is Woodward’s work [42], which demonstrated that infants distinguish between animate and inanimate entities in dynamic scenes, and attribute goals to agents or animate actors, but not to inanimate objects. This finding has been validated and elaborated through other experiments and variations [43, 44, 2]. A study by Luo and Baillargeon [27] extended the understanding of goal attribution by showing that infants can attribute goals to new self-propelled objects. Further research illustrated that infants can generalize the attribution of goals to new actions when provided with sufficient contextual cues [18]. In addition, infants exhibit a significant understanding of objects and their properties from an early age [1, 37, 16], expecting objects in the scene to follow physical rules, unlike animate agents that follow their goals [36, 25].
2.2 Related vision models of infants early learning
A large number of empirical studies have shown impressive capacities of young infants to understand different aspects of agent-object interactions in a dynamic scene, including the ownership of an object [10, 33], Efficiency of agent-object interaction [12, 29], and using objects as tools [24]. In parallel to empirical studies of infant learning, computational models, including computer vision models, have been developed and used to replicate and better understand a range of related empirical findings.
@@ 2.2 Related vision models of infants early learning
A recent study [38] investigated goal attribution in infants, both empirically and in vision models. The study found that infants anticipate agents’ actions to be directed at objects rather than locations. In contrast, AI models often target locations, highlighting the need to integrate infants’ understanding into models to better replicate human behavior. Li’s work [23] introduced additional tasks related to agents and objects, along with a self-supervised model. This work demonstrated limitations of current neural network models in goal attribution tasks.
@@
Several works [3, 15, 45] have addressed infant-like tasks, using the Baby Intuitions Benchmark (BIB) [11], a dataset specifically designed for evaluating developmental tasks related to agents. These studies employed various approaches, such as transformer-based architectures and Bayesian models, to tackle these tasks.
@@ 3 Method
3.1 Goal-Directed Dataset
We designed a dataset inspired by Woodward’s experiments [42], extending it to include tasks beyond those studied in prior research. The dataset consists of sequences of simple scenes (Figure 1), created using icons from ’flaticon.com’. Each frame contains three entities: two inanimate objects and an actor, which can be either animate (e.g., hands, animals, people) or inanimate (e.g., books, flowers).
@@ 3.2 The Two Models
3.2.1 First Step - Scene Representations
The objective of the first step is to create scene representations for the image sequences. To achieve this, we fine-tuned the BLIP model [22] to generate representations for each frame, encoding the class and location of the three entities (upper-left, upper-right, bottom). For example, the representation for the first frame in the top row of Figure 1 would be: ’cake’, ’bag’, ’girl’. The meaning of the words (e.g., ’cake’, ’bag’, ’girl’) is not used by the model. Instead, the words only serve as placeholders to encode the entities’ identities.
@@ 3.2.1 First Step - Scene Representations
3.2.2 Second Step - Prediction Generation
The prediction task is framed as a classification problem using the BERT model [6]. The input to the model is the full sequence representation of the frames, and the model’s output is a prediction of the actor’s future location (’left’ or ’right’) in the next step of the sequence. The results in the following sections evaluate the model’s accuracy in this prediction. Training and optimization details are provided in Appendix A.1.
We used large models (BLIP, BERT) in our cognitive model to reflect the complexity of the infant brain, as infants, too, are born with a sophisticated neural architecture. We fine-tuned all layers in both models for our specific tasks. To ensure that the pre-trained linguistic knowledge of the models did not drive the observed effects, we also ran a version of the experiment using binary representations (e.g., 0/1 vectors) instead of words in natural language and obtained similar results. Furthermore, we confirmed that their prior knowledge did not bias the results by verifying that the initial performance of the models in our tasks was not significantly different from the chance level.
Figure 2: Two-step process of the cognitive and naive models: First, scene representations are created, and then predictions are made. Concepts specific only to the cognitive model are shown in bold and purple.
4 Experiments
4.1 Experiment 1: Prediction
@@
The primary question was how the two models compare in learning the prediction task. Two important points regarding this comparison are: 1. The naive model could theoretically learn to distinguish between the two types of actors based on the data it observes, as the actor type is consistently correlated with its behavior in both training and test scenarios. 2. Note that the addition, such as the labels ‘animate’ and ‘non-animate’, is provided in a ’bare’ form, without additional context about its implications or how it relates to the model’s predictions. For the model, these labels are arbitrary markers with no associated information.
4.1.1 Data
The models were trained using sequences of three frames (Figure 1), employing various combinations of icons, with half of the sequences showing animate actors and half showing non-animate actors. Two dataset sizes were used: a small dataset (320 training examples and 80 test examples) and a large dataset (four times larger, 1280 training, 320 test examples). The test data used the same actors as the training data, but introduced novel target objects to evaluate generalization. Both the cognitive and naive networks were trained for 2000 epochs. Each experiment was repeated 15 times with randomly generated sequences, and the results were averaged.
4.1.2 Results
The averaged results of Experiment 1 for both models are shown in Figure 3, comparing their performance on the small and large datasets. Only the first 1000 epochs are shown, as the results stabilized within this range. The cognitive network achieved perfect accuracy after 100 epochs with the smaller dataset, and converged to perfect accuracy after 50 epochs with the larger dataset. In contrast, the naive network achieved approximately 80% accuracy after 1000 epochs with the smaller dataset and 87% accuracy with the larger dataset.
@@ 4.1.2 Results
4.2 Experiment 2: Generalization
Generalization to new tasks and domains is a crucial aspect and a useful measure of intelligent behavior [20, 4]. In the previous experiment, the test data used the same actors as in the training but introduced new target objects. Here, we extended this to assess the networks’ ability to generalize by evaluating their performance with entirely new actors and a novel spatial configuration.
To address generalization, a 5-frame paradigm was introduced (Figure 4). This paradigm first presents the actor’s behavior in the initial three frames and then introduces the same actor with new target objects, requiring the model to predict the actor’s behavior in the next step. Unlike Experiment 1, this setup includes information about the actor’s behavior within the data itself, theoretically enabling both networks to predict an actor’s behavior from a single example. As a result, it is well-suited for testing generalization to new actors.
4.2.1 Experiment 2.1: Generalization to New Actors
Data. The 5-frame paradigm incorporates the second and third frames from Experiment 1 (Prediction task), along with the frame that features the next step of the sequence, which was not shown in Experiment 1. The first three frames of Experiment 2 already show the actor’s preference. This sequence is extended by adding two new frames that present the same actor with new target objects. The prediction task focuses on the actor’s behavior in relation to these new targets. The models were trained using sequences of five frames, each featuring various combinations of icons, with half of the sequences containing animate actors and the other half containing non-animate actors.
Two task types were generated: Task 1 (T1): A simpler task, where the test data uses the same actors as in the training data but introduces new target objects. Task 2 (T2): A more challenging task because the test data includes entirely new actors and new target objects.
The networks were first trained on T1 and then further fine-tuned under one of two conditions: T1-T1: The same simple task (T1) with additional data. This condition served as a control to assess the effect of additional data without introducing new challenges. T1-T2: Following the initial task T1, the model is fine-tuned for the more challenging generalization task (T2). Each task used 640 training examples and 160 test examples. Both networks were trained for 1000 epochs. Each experiment was repeated 15 times with randomly generated sequences, and the results were averaged.
@@ 4.2 Experiment 2: Generalization
Results. The results for both models are shown in Figure 5, comparing the performance under the following conditions: 1. Learning the initial task (T1). 2. Retraining on the same task with additional data (T1-T1). 3. Generalization to a new task (T1-T2). Only the first 300 epochs are shown, as there was no significant change afterward. The cognitive network outperformed the naive network in both accuracy and learning speed across all tasks: T1-T1 Condition: The cognitive network achieved near-perfect accuracy in the first run and 100% accuracy almost immediately in the second run. The naive network stabilized at 58% accuracy after 1000 epochs in the first run and reached 75% accuracy after 200 epochs in the second run. T1-T2 Condition: For the challenging generalization task, the cognitive network achieved perfect accuracy after just a few epochs. The naive network, however, reached only  65% accuracy, performing worse than in the T1-T1 condition.
@@
Figure 5: Experiment 2.1: Generalization to New Actors Results. Average test accuracy of the naive and cognitive networks on Task 1 (T1), retraining on Task 1 (T1-T1), and retraining on Task 2 (T1-T2), with SEM included.
4.2.2 Experiment 2.2: Spatial Generalization
Data. We used the same five-frame paradigm described above for the spatial generalization experiment. As in the previous section, both networks were first trained on Task 1 (T1). We then introduced two spatial generalization conditions. In the first condition (T1-Generalization), the networks were initially trained on T1 and then retrained on a spatially modified version of the task obtained by horizontally flipping the original five-frame configurations. For example, objects that previously appeared at the top of the frame were relocated to the bottom. The actors remained the same as in T1. In the second condition (T1-Spatial+Actors), the networks were trained on T1 and then retrained on the horizontally flipped configurations with entirely new actors. Each task used 640 training examples and 160 test examples. Both networks were trained for 300 epochs. Each experiment was repeated 15 times with randomly generated sequences, and the results were averaged.
Results. The results for both models are shown in Figures 6 and 7 for the two spatial generalization conditions, respectively. The initial T1 training phase is identical across both conditions (for clarity, we display the results from one representative run) and corresponds to the same T1 task used in Experiment 2.1. Minor differences in accuracy reflect natural variation across runs, as captured by the SEM. In this setting, the naive network reached accuracy of 61% after 100 epochs, while the cognitive network achieved near-perfect accuracy within the same number of epochs. In both spatial generalization conditions, the cognitive network outperformed the naive network in terms of both final accuracy and learning speed. In the T1-Generalization condition (horizontal spatial flip with the same actors), the cognitive network reached near-perfect accuracy on the second task within approximately 50 epochs. The naive network, in contrast, reached only 68% accuracy after 100 epochs and showed no substantial improvement thereafter. In the more challenging T1-Generalization+Actors condition (horizontal flip combined with new actors), the cognitive network again achieved near-perfect accuracy rapidly. The naive network reached approximately 63% accuracy after 150 epochs and maintained this level without further improvement.
Importantly, in both spatial generalization settings, the cognitive network began the second task with an initial accuracy higher than that at the beginning of the original T1 training. This suggests that knowledge acquired during T1 generalized to the modified tasks. In contrast, the naive network showed little evidence of such generalization: its initial performance on the spatially modified tasks was similar to its starting point in the original T1 phase, indicating minimal generalization from prior learning.
These results demonstrate that the difference between the cognitive and naive networks is not limited to final performance levels but also reflects differences in generalization abilities. The cognitive network not only achieves higher final accuracy, but also adapts more rapidly to spatially modified and actor-novel settings.
Figure 6: Experiment 2.2: Spatial Generalization Results. Average test accuracy of the naive and cognitive networks on Task 1 (T1) and retraining on Spatial task (T1-Spatial), with SEM included.
Figure 7: Experiment 2.2: Spatial Generalization with New Actors Results. Average test accuracy of the naive and cognitive networks on Task 1 (T1) and retraining on Spatial and New Actors task (T1-Spatial+Actors), with SEM included.
4.3 Experiment 3: Decomposition
We address the following question: How is the additional information provided to the cognitive model learned? In earlier experiments, the actor’s type (animate/inanimate) was explicitly provided to the cognitive network. In the current experiment, however, the same information is available to both models. Inspired by evidence that self-propelled motion is a strong indicator of animacy in infants [27], the experiment introduced additional frames to enable the learning of this concept.